OpenStudy (anonymous):

Variable y varies directly with variable x, and y = 18 when x = 4. What is the constant of variation in decimal form?

OpenStudy (igreen):

\(\sf y = kx\) Plug in what we know: \(\sf 18 = k(4)\) Divide 4 to both sides, what do you get?

OpenStudy (anonymous):

The constant of variation is 18/4 = 9/2= 4.5

OpenStudy (anonymous):

I got 4.5

OpenStudy (igreen):

4.5 is correct.

OpenStudy (anonymous):

Thanks :)